### v2

@@ 3.2 Inadequacy of Existing Approaches
However, even this sophisticated gradient computation does not resolve the long edge problem. The gradient formula incorporates edge masses that quantify neighborhood overlaps, but **the fundamental issue remains**: a long edge connecting vertices in different basins can still produce a gradient that directs flow across basin boundaries, even when the edge has appropriately small mass. The problem is not that the gradient is computed incorrectly, but that the graph topology itself contains edges that violate the manifold geometry. No local gradient computation can distinguish between a long edge that shortcuts across a valley (problematic) versus one that legitimately connects distant points along a genuine gradient path (acceptable).
@@ 4. Co-Monotonicity Coefficients
Having established gradient flow partitions through validated basins, we turn to quantifying associations between functions on graph vertices. The fundamental question driving this development is: how should we measure whether two functions vary together or oppositely along the edges of a graph? Classical correlation coefficients treat all observations as equally related, ignoring the geometric structure encoded by the graph. We seek measures that respect this structure, detecting directional concordance locally while adapting to spatial heterogeneity in association patterns.
4.1 The Inadequacy of Global Association Measures
@@
where $\bar{y}$ and $\bar{z}$ denote sample means. This measure treats all vertex pairs equally, implicitly assuming that observations at vertices $v$ and $u$ are as related as observations at $v$ and $w$, regardless of whether edges connect these pairs in the graph. When the graph encodes meaningful proximity relationships, as in spatial data or network analysis, this assumption discards valuable structural information.
Moreover, correlation measures linear association in terms of deviations from means, which may not align with the geometric properties of interest. Within a gradient flow cell where the outcome $\hat{y}$ varies monotonically, we seek features that exhibit monotone co-variation along the same paths, increasing when $\hat{y}$ increases and decreasing when $\hat{y}$ decreases. A feature $z$ that oscillates around a constant mean but always changes in the same direction as $\hat{y}$ across edges would receive zero Pearson correlation despite perfect directional concordance.
Mutual information and related entropy-based measures address nonlinearity and can be computed on graph structures, providing viable alternatives for quantifying associations between vertex functions. However, these measures produce a single global summary of the relationship between $y$ and $z$ across all vertices. They cannot reveal spatial heterogeneity in association patterns, nor can they identify regions where the relationship is strong versus weak, or positive versus negative. When association structure varies across the graph, as commonly occurs in gradient flow cells with different monotonic behaviors, a single global measure obscures this regional variation.
@@
4.2 Edge-Based Directional Concordance
We begin by formalizing the notion of directional agreement along a single edge. Let $e=[i,j]\in E$ be an edge connecting vertices $i$ and $j$, and define the edge difference operator acting on a function $f:V\to\mathbb{R}$ by
$\Delta_{e}f=f(j)-f(i).$ (12)
This operator measures the change in $f$ when moving from $i$ to $j$ along edge $e$. The orientation of edges is arbitrary; what matters is that we consistently use the same orientation for all functions when computing products of differences.
@@ 4.2 Edge-Based Directional Concordance
4.3 Vertex-Level Co-Monotonicity Coefficients
We aggregate edge-wise directional concordance over neighborhoods to obtain vertex-level measures. At each vertex $v$, let $N(v)$ denote the set of neighbors (vertices connected to $v$ by edges). We sum the products $\Delta_{e}y\cdot\Delta_{e}z$ over all edges incident to $v$, weighting by edge-specific factors $w_{e}\geq 0$, and normalize by the total weighted absolute co-variation.
@@ 4.3 Vertex-Level Co-Monotonicity Coefficients
Given functions $y,z:V\to\mathbb{R}$ on a weighted graph $G=(V,E)$ with edge weights $w_{e}\geq 0$, the co-monotonicity coefficient at vertex $v$ is
$\text{cm}(y,z;w)(v)=\frac{\sum_{u\in N(v)}w_{e}\Delta_{e}y\cdot\Delta_{e}z}{\sum_{u\in N(v)}w_{e}|\Delta_{e}y\cdot\Delta_{e}z|},$ (13)
where $e=[v,u]$ is the edge connecting $v$ and $u$, and $\Delta_{e}y=y(u)-y(v)$, $\Delta_{e}z=z(u)-z(v)$. When the denominator vanishes (all products $\Delta_{e}y\cdot\Delta_{e}z=0$), we define $\text{cm}(y,z;w)(v)=0$.
The normalization ensures that $\text{cm}(y,z;w)(v)\in[-1,1]$ for all vertices. Perfect positive co-monotonicity (both functions always change in the same direction within the neighborhood) yields $\text{cm}(y,z;w)(v)=1$. Perfect negative co-monotonicity (functions always change in opposite directions) yields $\text{cm}(y,z;w)(v)=-1$. Values near zero indicate no systematic directional relationship, either because the functions vary independently or because positive and negative concordance balance within the neighborhood.
4.4 Weighting Schemes and Geometric Interpretations
The edge weights $w_{e}$ control how we aggregate directional concordance over neighborhoods. Three canonical weighting schemes address different geometric and statistical contexts.
The simplest choice assigns uniform weight to all edges: $w_{e}=1$ for all $e\in E$. This unit weighting scheme treats edges equally regardless of their length, emphasizing the combinatorial structure of the graph. The coefficient $\text{cm}(y,z;1)(v)$ counts the proportion of edges where functions agree in direction, providing a robust summary insensitive to the magnitude of changes. Unit weighting is appropriate when edge lengths are roughly comparable, as in regular lattices, or when we wish to count directional agreements without geometric normalization. For graphs constructed from irregular spatial samples where edge lengths vary substantially, unit weighting may overweight long edges that span large distances in the underlying space.
To account for geometric scaling, we introduce derivative weighting. Let $\Delta_{e}$ denote the length (or weight) of edge $e$ in the graph metric. Setting $w_{e}=1/\Delta_{e}^{2}$ normalizes the coefficient to compare directional derivatives rather than absolute changes:
$\text{cm}_{\partial}(y,z)(v)=\frac{\sum_{u\in N(v)}\frac{1}{\Delta_{e}^{2}}\Delta_{e}y\cdot\Delta_{e}z}{\sum_{u\in N(v)}\frac{1}{\Delta_{e}^{2}}|\Delta_{e}y\cdot\Delta_{e}z|}=\frac{\sum_{u\in N(v)}\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}}{\sum_{u\in N(v)}\left|\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}\right|}.$ (14)
The quantities $\Delta_{e}y/\Delta_{e}$ and $\Delta_{e}z/\Delta_{e}$ approximate directional derivatives along edge $e$, making $\text{cm}_{\partial}(y,z)(v)$ analogous to measuring correlation between gradients in the continuous setting. Derivative weighting naturally mitigates the influence of long edges: a large change $\Delta_{e}y$ across a long edge $\Delta_{e}$ corresponds to a small derivative $\Delta_{e}y/\Delta_{e}$, reducing its contribution to the coefficient. This weighting is appropriate when functions represent continuous quantities sampled at irregular positions, and we wish to assess co-variation in the underlying continuous fields.
For robustness to outliers and heavy-tailed distributions, sign-based weighting uses only the direction of change, discarding magnitude information entirely. We modify the definition slightly, replacing weighted sums with counts:
$\text{cm}_{\pm}(y,z)(v)=\frac{\sum_{u\in N(v)}\text{sign}(\Delta_{e}y\cdot\Delta_{e}z)}{|N(v)|},$ (15)
where $\text{sign}(x)=1$ if $x>0$, $\text{sign}(x)=-1$ if $x<0$, and $\text{sign}(0)=0$. This coefficient measures the proportion of edges where functions agree in direction minus the proportion where they disagree. Sign-based weighting is insensitive to the magnitude of function changes, making it robust to outliers or to situations where only ordinal relationships are meaningful. It trades geometric precision for statistical robustness.
4.5 Alternative Normalization and Geometric Interpretation
The co-monotonicity coefficient defined in Section 4.3 uses absolute-value normalization to measure the proportion of concordant versus discordant directional changes across edges. We now introduce a complementary measure with Pearson-style normalization that weights edges by the magnitudes of their changes. This variant admits a rigorous geometric interpretation as the cosine of the angle between gradient vectors on smooth manifolds, providing theoretical justification for the derivative-weighted form. The two normalizations capture related but distinct aspects of directional concordance, and their differences prove informative for understanding local association structure.
Correlation-Type Co-Monotonicity Coefficient
We begin with the alternative definition. Given functions $y,z:V\to\mathbb{R}$ on a weighted graph $G=(V,E)$ with edge weights $w_{e}\geq 0$, the correlation-type co-monotonicity coefficient at vertex $v$ is
@@ 4.5 Alternative Normalization and Geometric Interpretation
where $e=[v,u]$ denotes the edge connecting $v$ and $u$, and $\Delta_{e}y=y(u)-y(v)$, $\Delta_{e}z=z(u)-z(v)$ are the edge differences. When either denominator term vanishes (indicating that one of the functions is constant across all edges incident to $v$), we define $\text{cm}_{\text{cor}}(y,z;w)(v)=0$ by convention.
The normalization by geometric mean of squared changes ensures that $\text{cm}_{\text{cor}}(y,z;w)(v)\in[-1,1]$ for all vertices, with extremal values achieved under perfect proportionality. If $\Delta_{e}z=k\Delta_{e}y$ for some constant $k>0$ across all edges incident to $v$, then $\text{cm}_{\text{cor}}(y,z;w)(v)=1$. If $\Delta_{e}z=-k\Delta_{e}y$ for $k>0$, then $\text{cm}_{\text{cor}}(y,z;w)(v)=-1$. The measure vanishes when positive and negative products balance in a squared-magnitude-weighted sense, which differs subtly from the absolute-value-weighted balance in the definition of Section 4.3.
For derivative-weighted co-monotonicity where $w_{e}=1/\Delta_{e}$ with $\Delta_{e}$ denoting edge length, the coefficient takes the form
$\text{cm}_{\text{cor},\partial}(y,z)(v)=\frac{\sum_{u\in N(v)}\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}}{\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}y}{\Delta_{e}}\right)^{2}}\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}z}{\Delta_{e}}\right)^{2}}},$ (17)
which directly correlates the discrete directional derivatives across the neighborhood of $v$. This form proves essential for the manifold interpretation we develop below.
Geometric Interpretation via Gradient Alignment

### v3

@@ 3.2 Inadequacy of Existing Approaches
However, even this sophisticated gradient computation does not resolve the long edge problem. The gradient formula incorporates edge masses that quantify neighborhood overlaps, but the fundamental issue remains: a long edge connecting vertices in different basins can still produce a gradient that directs flow across basin boundaries, even when the edge has appropriately small mass. The problem is not that the gradient is computed incorrectly, but that the graph topology itself contains edges that violate the manifold geometry. No local gradient computation can distinguish between a long edge that shortcuts across a valley (problematic) versus one that legitimately connects distant points along a genuine gradient path (acceptable).
@@ 4. Co-Monotonicity Coefficients
Having established gradient flow partitions through local extrema gradient flow basins, we turn to quantifying associations between functions on graph vertices. The fundamental question driving this development is: how should we measure whether two functions vary together or oppositely along the edges of a graph? Classical correlation coefficients treat all observations as equally related, ignoring the geometric structure encoded by the graph. We seek measures that respect this structure, detecting directional concordance locally while adapting to spatial heterogeneity in association patterns.
4.1 Global Association Measures
@@
where $\bar{y}$ and $\bar{z}$ denote sample means. This measure treats all vertex pairs equally, implicitly assuming that observations at vertices $v$ and $u$ are as related as observations at $v$ and $w$, regardless of whether edges connect these pairs in the graph. When the graph encodes meaningful proximity relationships, as in spatial data or network analysis, this assumption discards valuable structural information. In particular, when a feature $z$ varies concordantly with $y$ in one region of the graph but discordantly in another, these opposing local relationships may cancel in the global correlation, yielding a coefficient near zero despite strong region-specific associations.
Moreover, correlation measures linear association in terms of deviations from means, which may not align with the geometric properties of interest. Within a gradient flow cell where the outcome $y$ varies monotonically, we seek features that exhibit monotone co-variation along the same paths, increasing when $y$ increases and decreasing when $y$ decreases. Mutual information addresses nonlinearity and can be computed on graph structures, providing a viable alternative for quantifying associations between vertex functions. However, mutual information and other global measures produce a single summary of the relationship between $y$ and $z$ across all vertices. They cannot reveal spatial heterogeneity in association patterns, nor can they identify regions where the relationship is strong versus weak, or positive versus negative. When association structure varies across the graph, as commonly occurs in gradient flow cells with different monotonic behaviors, a single global measure obscures this regional variation. While mutual information can in principle be localized to vertices or paths through kernel-weighted density estimation (see Appendix B), such approaches introduce substantial practical challenges in estimation and interpretation.
@@
We formalize these concepts through two complementary perspectives on directional concordance. At each vertex $v\in V$, we define a coefficient $c(y,z;v)\in[-1,1]$ that measures whether $y$ and $z$ tend to increase or decrease together across edges incident to $v$. This local measure naturally extends to paths $\gamma=(\gamma_{0},\gamma_{1},\ldots,\gamma_{n})$, where we compute $c(y,z;\gamma)$ to quantify directional concordance along the sequence of edges comprising $\gamma$. The vertex-based coefficients enable spatial mapping of association patterns, while path-based coefficients characterize monotonic behavior along gradient flow trajectories. We develop both perspectives in the following sections, beginning with the vertex-centered formulation.
4.2 Vertex-Level Co-Monotonicity Coefficients
We seek vertex-level measures that quantify whether two functions increase or decrease together across edges in local neighborhoods. The approach builds on edge-wise directional concordance, aggregating signed products of function changes across incident edges with appropriate normalization to yield coefficients bounded in [1,1][-1,1].
We begin by formalizing the notion of directional agreement along a single edge. Let $e=[v,u]\in E$ be an edge connecting vertices $v$ and $u$, and define the edge difference operator acting on a function $f:V\to\mathbb{R}$ by
$\Delta_{e}f=f(u)-f(v).$ (12)
This operator measures the change in $f$ when moving from $v$ to $u$ along edge $e$. The choice of orientation (which vertex to designate as initial point) is arbitrary; what matters is that we consistently use the same orientation for all functions when computing products of differences, ensuring that the product captures directional concordance rather than being affected by orientation choices.
@@ 4.2 Vertex-Level Co-Monotonicity Coefficients
To construct a vertex-level measure, we aggregate these products over edges incident to each vertex. Let $N(v)$ denote the set of neighbors of $v$ (vertices connected to $v$ by edges). A natural approach sums the weighted products $w_{e}\Delta_{e}y\cdot\Delta_{e}z$ over all edges $e=[v,u]$ with $u\in N(v)$, where $w_{e}\geq 0$ are edge weights. However, this raw sum depends on the scales of $y$ and $z$, making comparisons across different function pairs problematic. We address this by normalizing in a manner analogous to Pearson correlation, dividing by the geometric mean of the total weighted squared changes in each function separately. This normalization ensures the resulting coefficient lies in [1,1][-1,1] regardless of function scales, with extremal values indicating perfect proportional co-variation or anti-variation.
@@
Given functions $y,z:V\to\mathbb{R}$ on a weighted graph $G=(V,E)$ with edge weights $w_{e}\geq 0$, the correlation-type co-monotonicity coefficient at vertex $v$ is
@@
where $e=[v,u]$ denotes the edge connecting $v$ and $u$, and $\Delta_{e}y=y(u)-y(v)$, $\Delta_{e}z=z(u)-z(v)$ are the edge differences oriented away from $v$. When either denominator term vanishes (indicating that one of the functions is constant across all edges incident to $v$), we define $\text{cm}_{\text{cor}}(y,z;w)(v)=0$ by convention.
The normalization by geometric mean of squared changes ensures that $\text{cm}_{\text{cor}}(y,z;w)(v)\in[-1,1]$ for all vertices, with extremal values achieved under perfect proportionality. If $\Delta_{e}z=k\Delta_{e}y$ for some constant $k>0$ across all edges incident to $v$, then $\text{cm}_{\text{cor}}(y,z;w)(v)=1$. If $\Delta_{e}z=-k\Delta_{e}y$ for $k>0$, then $\text{cm}_{\text{cor}}(y,z;w)(v)=-1$. The coefficient vanishes when positive and negative products balance in a squared-magnitude-weighted sense.
A particularly important special case arises when edge weights are chosen as $w_{e}=1/\ell_{e}$ where $\ell_{e}$ denotes the length of edge $e$. This derivative-weighted co-monotonicity coefficient takes the form
$\text{cm}_{\text{cor},\partial}(y,z)(v)=\frac{\sum_{u\in N(v)}\frac{\Delta_{e}y}{\ell_{e}}\cdot\frac{\Delta_{e}z}{\ell_{e}}}{\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}y}{\ell_{e}}\right)^{2}}\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}z}{\ell_{e}}\right)^{2}}},$ (14)
where the ratios $\Delta_{e}y/\ell_{e}$ and $\Delta_{e}z/\ell_{e}$ approximate directional derivatives of $y$ and $z$ along the edges. This form admits a compelling geometric interpretation: on a smooth Riemannian manifold, the coefficient measures the cosine of the angle between gradient vector fields, providing theoretical justification for this normalization choice. We develop this connection formally in Section 4.4, where we show that $\text{cm}_{\text{cor},\partial}$ arises naturally from the Riemannian metric structure on smooth manifolds.
Throughout the remainder of this paper, we adopt the simplified notation $\text{cm}(y,z;w)(v)$ to denote the correlation-type co-monotonicity coefficient $\text{cm}_{\text{cor}}(y,z;w)(v)$ when the normalization scheme is clear from context. For derivative weighting, we write $\text{cm}_{\partial}(y,z)(v)$ as shorthand for $\text{cm}_{\text{cor},\partial}(y,z)(v)$. The subscript ”cor” is retained only when explicitly contrasting with alternative normalizations.
4.3 Geometric Interpretation via Gradient Alignment